OpenStudy (anonymous):

so i used the formula sn=n/2[2a+(n−1)d] but got -435 and that's not the right answer, someone help me find out what i did wrong? Find S15 for the series -1 + -3 + -5 + -7 +…

OpenStudy (anonymous):

can u find d??

OpenStudy (anonymous):

Common difference

OpenStudy (anonymous):

Can you show your work ??

OpenStudy (anonymous):

d = t2 -t1

OpenStudy (anonymous):

d=t2-t1

OpenStudy (anonymous):

@waterineyes i think we should have went wrong in finding d

OpenStudy (anonymous):

that is the only chance or calcu mistake

OpenStudy (anonymous):

@Yahoo! let her give what she did..

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

@summer1nthec1ty we can find your mistake can you show your work how you did it ??

OpenStudy (anonymous):

[15/2(2(-1)+(15-1)*-4=-435

OpenStudy (anonymous):

aha.......... d is wrong

OpenStudy (anonymous):

d = t2 - t1

OpenStudy (anonymous):

Difference is wrong I think...

OpenStudy (anonymous):

Dang itt

OpenStudy (anonymous):

-3 - (-1 ) = -2

OpenStudy (anonymous):

\[\large d = -3 - (-1) = ??\]

OpenStudy (anonymous):

-(-) = +

OpenStudy (anonymous):

SHOOT thanks guys

OpenStudy (anonymous):

welcome..))